# Inferring parameters and reconstruction of two-dimensional turbulent flows with physics-informed neural networks

### v1

@@ Abstract
Extracting information and reconstructing a flow field from sparse or noisy measurements is important yet difficult in many applications. Solving the problem requires additional knowledge to compensate for imperfections in the data. A well-constructed physical model of the phenomenon can serve in this capacity. However, traditional computational fluid dynamics (CFD) models are ill-suited for these purposes due to the computational complexity of data assimilation methods [1], especially for turbulent flows. Recent advances in machine learning (ML) open up new opportunities for addressing these challenges [2, 3, 4, 5].
In this work, we will focus on physics-informed neural networks (PINNs) [6], which provide a framework for merging data with physical laws expressed as partial differential equations (PDEs). Previously, PINNs were applied to uncover near-wall blood flow from sparse data [7], for super-resolution and denoising applications [8, 9, 10], to reconstruct velocity and pressure fields from flow visualizations [11, 12] and particle image/tracking velocimetry (PIV/PTV) data [13, 14], to simulate turbulence [15, 16, 17], just to name a few. A more comprehensive bibliography can be found in recent reviews [3, 4, 18, 5]. Here we implement PINN to solve inverse problem inspired by experimental studies of forced turbulent flows in thin conducting fluid layers [19, 20, 21, 22, 23, 24]. Our goal is to show that when the physical model is taken into account, the suggested technique has the ability to extract more information from conventional flow observations.
Let us start by defining the problem. We consider two-dimensional incompressible fluid flow described by the forced Navier-Stokes equation with linear bottom friction
@@
where $\bm{v}$ is 2D velocity, $p$ is the pressure, $\alpha$ is the friction coefficient, and $\nu$ is the kinematic viscosity. The external force $\bm{f}$ is assumed to be stationary in time and divergence-free in space. We are interested in a moderately turbulent regime, when the nonlinear term in the Navier-Stokes equation plays a substantial role. Then, despite the stationary nature of the external forcing, the fluid flow is characterized by strong spatiotemporal variations. Now suppose that someone measures the velocity field in a certain region (possibly with open boundaries) and makes observations for several times longer than the typical time of flow fluctuations. In what follows, we will demonstrate how, using ML techniques, one can not only reconstruct the dense velocity $\bm{v}(\bm{r},t)$ and pressure $p(\bm{r},t)$ fields (super-resolution application), but also establish the driving force $\bm{f}(\bm{r})$ and unknown fluid viscosity $\nu$ and bottom friction $\alpha$, based only on potentially sparse and noisy velocity field data. Let us emphasize that the developed approach does not require knowledge of the boundary conditions at the bounds of the observation area.
Data generation. Data for this study are obtained by integrating (1) using the GeophysicalFlows.jl pseudospectral code [25], which has recently been successfully applied to model two-dimensional turbulence [26, 27, 28]. The code can be executed on the GPU, resulting in high computational performance. The aliasing errors are removed with the two-third rule. The simulation domain is a doubly periodic box of size $2\pi\times 2\pi$. The boundary conditions and form of the domain are only relevant to the data production process; the rest of the study is generic in this regard.
@@
Fig. 1a shows snapshot of the vorticity field $\omega=\partial_{x}v_{y}-\partial_{y}v_{x}$ in the statistical steady-state. The flow is chaotic, with no obvious imprint of an external forcing. Since there are two dissipation mechanisms – viscosity and bottom friction – the flow state can be characterized by two dimensionless parameters [29]: the Reynolds number $Re=UL/\nu\approx 1.3\times 10^{3}$ and $Rh=U/(\alpha L)\approx 13$, which were computed using the root mean square velocity $U$ and the size $L$ of the observation domain. The energy spectrum is presented in Fig. 1b. It indicates that fluid flow is mainly determined by modes with wave vectors $k\lesssim k_{f}$.
@@
Next, we assume that measurements reveal the velocity field $\bm{v}^{d}(\bm{r}_{n}^{d},t_{n}^{d})$ at $N_{data}$ points $\{\bm{r}_{n}^{d},t_{n}^{d}\}_{n=1}^{N_{data}}$, which are randomly scattered within the observation area and time interval of length $T$. These data imitate PIV/PTV measurements and will be used for subsequent neural network training. To study the influence of measurement noise, we will distort the velocity as $v_{x,y}^{d}\rightarrow(1+\eta)v_{x,y}^{d}$, where $\eta\sim\mathcal{N}(0,\varepsilon^{2})$ is a normally distributed random variable and the standard deviation $\varepsilon$ controls the noise level. We set $N_{data}=3\times 10^{4}$, which is around $0.2\%$ of the total data and thus represents sparse measurements.
Physics-Informed Neural Network. In our study, we utilize two independent neural networks (NNs) that are trained simultaneously, see Fig. 2a. The first NN approximates the velocity and pressure at a particular position and time. It has three input neurons and three output neurons, $(v_{x},v_{y},p)=\mathcal{F}_{\vec{\theta}_{1}}(x,y,t)$. The second NN is intended to predict stationary force. As a result, it contains only two input neurons and two output neurons, $(f_{x},f_{y})=\mathcal{G}_{\vec{\theta}_{2}}(x,y)$. The parameters $\vec{\theta}_{1}$ and $\vec{\theta}_{2}$ denote the trainable variables for NNs.
@@
NNs must be sufficiently complicated to reproduce patterns that are essential for the considered turbulent flow. Universal approximation theorem ensures that this is possible [30, 31]. In the original paper [6], the authors used a fully connected NN architecture, but here we implement a residual design [32] for both NNs as we found that it performs better in agreement with the results reported earlier [33]. The adopted architecture is shown in Fig. 2b. The parameters $N_{l}$ and $N_{n}$ represent the number of hidden layers and neurons in each hidden layer, respectively. The output $\bm{Y}_{k}$ of $k$-layer is given by
@@
where $\bm{Y}_{-1}\equiv\bm{0}$ for convenience of notation, $\bm{Y}_{0}$ corresponds to the input layer, $\bm{Y}_{N_{l}+1}$ denotes the output layer, $\hat{W}$ and $\bm{b}$ are weights and biases (trainable parameters that comprise $\vec{\theta}$), and $\sigma$ is the nonlinear activation function applied element-wise. We use a tanh activation function because it gives a reasonable trade-off between training time and resulting accuracy [13].
In general, increasing the size of NNs can reduce the prediction error, but only when the NNs are given enough data [13]. Here we set $N_{l}=7$ and $N_{n}=250$ for $\mathcal{F}_{\vec{\theta}_{1}}(\bm{r},t)$, and $N_{l}=5$ and $N_{n}=30$ for $\mathcal{G}_{\vec{\theta}_{2}}(\bm{r})$. This choice of parameters is consistent with $N_{data}=3\times 10^{4}$. Note that $\mathcal{G}_{\vec{\theta}_{2}}(\bm{r})$ is much simpler than $\mathcal{F}_{\vec{\theta}_{1}}(\bm{r},t)$ because it is not time dependent and the external forcing usually does not have a complex spatial dependency.
@@
Computing the right-hand sides involves differentiating the outputs of NNs with respect to input neurons using automatic differentiation [34]. The number and location of collocation points are not limited by measurements and can be chosen arbitrarily. We set $N_{eq}=3N_{data}$, with one-third of these points corresponding to data points $\{\bm{r}_{n}^{d},t_{n}^{d}\}_{n=1}^{N_{data}}$, and the rest are generated randomly. Note that instead of requiring the incompressibility conditions to be satisfied, the NNs can be modified to predict the relevant stream functions [6, 16]. However, in this case, calculating the terms in the Navier-Stokes equation involves computation of higher-order derivatives, which significantly slows down the training process.
@@
where $\beta$ is a weighting coefficient used to balance two terms. Its value can be adaptively changed during training [15], but here we use a simpler strategy in which the weight is fixed. Empirically, we have found that $\beta=0.02$ is well suited for this type of problem.
Minimization of the loss function (9) is performed by a gradient-based optimizer that adjusts network and model parameters. As the bottom friction and fluid viscosity are positive, it is convenient to represent them in the form $\alpha=e^{\lambda_{1}}$ and $\nu=e^{\lambda_{2}}$. So, the training variables are $\{\vec{\theta}_{1},\vec{\theta}_{2},\lambda_{1},\lambda_{2}\}$. First, we train the model for $10^{5}$ epochs using Adam optimizer [35] with learning rate of $\mu=10^{-3}$. To reduce computation time, we have implemented a mixed precision technique [36]. Then we use L-BFGS-B optimizer [37] until the tolerance reaches machine precision. The training is carried out on an NVIDIA A100 card. The code is written using the PyTorch library [38] and is openly available at [39].
@@
In conclusion, we demonstrated that PINN can integrate measurement data with an accurate physical model to extract additional information from ordinary flow observations. The developed approach was based on joint training of two independent NNs, each of which took into account the properties of individual terms of the Navier-Stokes equation, encoding them in its own architecture. The results obtained for turbulent flow showed good accuracy for super-resolution and inference applications and demonstrated moderate robustness to measurement noise. The suggested technique may be useful for analyzing observations (e.g., of the ocean or atmosphere), when it is difficult to carry out additional studies aimed at establishing the parameters (e.g., turbulent viscosity) of the physical model.

### v2

# Enhancing capabilities of particle image/tracking velocimetry with physics-informed neural networks

@@ Abstract
I Introduction
Extracting information and reconstructing a flow field from sparse or noisy measurements is important yet difficult in many applications. Solving the problem requires additional knowledge to compensate for imperfections in the data. A well-constructed physical model of the phenomenon can serve in this capacity. However, traditional computational fluid dynamics (CFD) models are ill-suited for these purposes due to the computational complexity of data assimilation methods Asch, Bocquet, and Nodet (2016), especially for turbulent flows. Recent advances in machine learning (ML) open up new opportunities for addressing these challenges Brunton, Noack, and Koumoutsakos (2020); Cai et al. (2021a); Karniadakis et al. (2021); Sharma et al. (2023).
In this work, we focus on physics-informed neural networks (PINNs) Lagaris, Likas, and Fotiadis (1998); Raissi, Perdikaris, and Karniadakis (2019), which provide a framework for merging data with physical laws expressed as partial differential equations (PDEs). Previously, PINNs were applied to uncover near-wall blood flow from sparse data Arzani, Wang, and D’Souza (2021), for super-resolution and denoising applications Fathi et al. (2020); Gao, Sun, and Wang (2021); Eivazi and Vinuesa (2022), to reconstruct velocity and pressure fields from flow visualizations Raissi, Yazdani, and Karniadakis (2020); Cai et al. (2021b); Clark Di Leoni et al. (2023a) and particle image/tracking velocimetry (PIV/PTV) data Wang, Liu, and Wang (2022); Clark Di Leoni et al. (2023b), to simulate turbulence Jin et al. (2021); Kag, Seshasayanan, and Gopinath (2022); Eivazi et al. (2022), just to name a few. A more comprehensive bibliography can be found in recent reviews Cai et al. (2021a); Karniadakis et al. (2021); Cuomo et al. (2022); Sharma et al. (2023). Here we implement PINN to solve inverse problem inspired by experimental studies of forced turbulent flows in thin conducting fluid layers Sommeria (1986); Paret and Tabeling (1998); Boffetta et al. (2005); Xia, Shats, and Falkovich (2009); Orlov, Brazhnikov, and Levchenko (2018); Fang and Ouellette (2021). Our goal is to show that when the physical model is taken into account, the suggested technique has the ability to extract more information from conventional flow observations.
We consider two-dimensional incompressible fluid flow described by the forced Navier-Stokes equation with linear bottom friction
@@ I Introduction
where $\bm{v}$ is 2D velocity, $p$ is the pressure, $\alpha$ is the friction coefficient, and $\nu$ is the kinematic viscosity. The external force $\bm{f}$ is assumed to be stationary in time and divergence-free in space. We are interested in a moderately turbulent regime, when the nonlinear term in the Navier-Stokes equation plays a substantial role. Then, despite the stationary nature of the external forcing, the fluid flow is characterized by strong spatiotemporal variations. Now suppose that someone measures the velocity field in a certain region (possibly with open boundaries) and makes observations for several times longer than the typical time of flow fluctuations. In what follows, we will demonstrate how to reconstruct the dense velocity $\bm{v}(\bm{r},t)$ and pressure $p(\bm{r},t)$ fields, and establish the driving force $\bm{f}(\bm{r})$ and unknown fluid viscosity $\nu$ and bottom friction $\alpha$, based only on potentially sparse and noisy velocity field data. Let us emphasize that the developed approach can be easily implemented on top of the existing PIV/PTV workflows, thereby expanding their functionality.
The contribution of this paper compared to previous studies is twofold. First, we generate a novel dataset of forced two-dimensional fluid flow that can serve as a benchmark solution in future research Hao et al. (2023). Second, we extend the PINN algorithm by including the ability to infer driving forces. This is achieved by simultaneously training two neural networks (NNs), one for the velocity and pressure field and the other for the external force. This division, in particular, makes it possible to explicitly take into account the stationarity of the external forcing in the design of the corresponding NN. Including physical information in NN architecture is known to enhance training performance and prediction accuracy.
II Data generation
Data for this study are obtained by integrating (1) using the GeophysicalFlows.jl pseudospectral code Constantinou et al. (2021), which has recently been successfully applied to model two-dimensional turbulence Parfenyev (2022); Kolokolov, Lebedev, and Parfenyev (2024); Parfenyev (2024). The code can be executed on the GPU, resulting in high computational performance. The aliasing errors are removed with the two-third rule. The simulation domain is a doubly periodic box of size $2\pi\times 2\pi$. The boundary conditions and form of the domain are only relevant to the data production process; the rest of the study is generic in this regard.
@@ II Data generation
Fig. 1a shows snapshot of the vorticity field $\omega=\partial_{x}v_{y}-\partial_{y}v_{x}$ in the statistical steady-state. The flow is chaotic, with no obvious imprint of an external forcing. Since there are two dissipation mechanisms – viscosity and bottom friction – the flow state can be characterized by two dimensionless parameters Mishra et al. (2015): the Reynolds number $Re=UL/\nu\approx 1.3\times 10^{3}$ and $Rh=U/(\alpha L)\approx 13$, which were computed using the root mean square velocity $U$ and the size $L$ of the observation domain. The energy spectrum is presented in Fig. 1b. It indicates that fluid flow is mainly determined by modes with wave vectors $k\lesssim k_{f}$.
@@
Next, we assume that measurements reveal the velocity field $\bm{v}^{d}(\bm{r}_{n}^{d},t_{n}^{d})$ at $N_{data}$ points $\{\bm{r}_{n}^{d},t_{n}^{d}\}_{n=1}^{N_{data}}$, which are randomly scattered within the observation area and time interval of length $T$. These data imitate PIV/PTV measurements and will be used for subsequent neural network training. To study the influence of measurement noise, we will distort the velocity as $v_{x,y}^{d}\rightarrow(1+\eta)v_{x,y}^{d}$, where $\eta\sim\mathcal{N}(0,\varepsilon^{2})$ is a normally distributed random variable and the standard deviation $\varepsilon$ controls the noise level. We set $N_{data}=3\times 10^{4}$, which is around $0.2\%$ of the total data and thus represents sparse measurements. On average, this corresponds to $150$ points per snapshot compared to $65536$ points at full resolution.
III Physics-Informed Neural Network
In our study, we utilize two independent NNs that are trained simultaneously, see Fig. 2a. The first NN approximates the velocity and pressure at a particular position and time. It has three input neurons and three output neurons, $(v_{x},v_{y},p)=\mathcal{F}_{\vec{\theta}_{1}}(x,y,t)$. The second NN is intended to predict stationary force. As a result, it contains only two input neurons and two output neurons, $(f_{x},f_{y})=\mathcal{G}_{\vec{\theta}_{2}}(x,y)$. The parameters $\vec{\theta}_{1}$ and $\vec{\theta}_{2}$ denote the trainable variables for NNs.
@@ III Physics-Informed Neural Network
NNs must be sufficiently complicated to reproduce patterns that are essential for the considered turbulent flow. Universal approximation theorem ensures that this is possible Cybenko (1989); Hornik, Stinchcombe, and White (1989). In the original paper Raissi, Perdikaris, and Karniadakis (2019), the authors used a fully connected NN architecture, but here we implement a residual design He et al. (2016) for both NNs as we found that it performs better in agreement with the results reported earlier Cheng and Zhang (2021). The adopted architecture is shown in Fig. 2b. The parameters $N_{l}$ and $N_{n}$ represent the number of hidden layers and neurons in each hidden layer, respectively. The output $\bm{Y}_{k}$ of $k$-layer is given by
@@
where $\bm{Y}_{-1}\equiv\bm{0}$ for convenience of notation, $\bm{Y}_{0}$ corresponds to the input layer, $\bm{Y}_{N_{l}+1}$ denotes the output layer, $\hat{W}$ and $\bm{b}$ are weights and biases (trainable parameters that comprise $\vec{\theta}$), and $\sigma$ is the nonlinear activation function applied element-wise. We use a tanh activation function because it gives a reasonable trade-off between training time and resulting accuracy Wang, Liu, and Wang (2022).
In general, increasing the size of NNs can reduce the prediction error, but only when the NNs are given enough data Wang, Liu, and Wang (2022). Here we set $N_{l}=7$ and $N_{n}=250$ for $\mathcal{F}_{\vec{\theta}_{1}}(\bm{r},t)$, and $N_{l}=5$ and $N_{n}=30$ for $\mathcal{G}_{\vec{\theta}_{2}}(\bm{r})$. This choice of parameters is consistent with $N_{data}=3\times 10^{4}$. Note that $\mathcal{G}_{\vec{\theta}_{2}}(\bm{r})$ is much simpler than $\mathcal{F}_{\vec{\theta}_{1}}(\bm{r},t)$ because it is not time dependent and the external forcing usually does not have a complex spatial dependency.
@@
Computing the right-hand sides involves differentiating the outputs of NNs with respect to input neurons using automatic differentiation Baydin et al. (2018). The number and location of collocation points are not limited by measurements and can be chosen arbitrarily. We set $N_{eq}=3N_{data}$, with one-third of these points corresponding to data points $\{\bm{r}_{n}^{d},t_{n}^{d}\}_{n=1}^{N_{data}}$, and the rest are generated randomly. Note that instead of requiring the incompressibility conditions to be satisfied, the NNs can be modified to predict the relevant stream functions Raissi, Perdikaris, and Karniadakis (2019); Kag, Seshasayanan, and Gopinath (2022). However, in this case, calculating the terms in the Navier-Stokes equation involves computation of higher-order derivatives, which significantly slows down the training process.
@@
where $\beta$ is a weighting coefficient used to balance two terms. Its value can be adaptively changed during training Wang, Teng, and Perdikaris (2021); Jin et al. (2021), but here we use a simpler strategy in which the weight is fixed. Empirically, we have found that $\beta=0.02$ is well suited for this type of problem.
Minimization of the loss function (9) is performed by a gradient-based optimizer that adjusts network and model parameters. As the bottom friction and fluid viscosity are positive, it is convenient to represent them in the form $\alpha=e^{\lambda_{1}}$ and $\nu=e^{\lambda_{2}}$. So, the training variables are $\{\vec{\theta}_{1},\vec{\theta}_{2},\lambda_{1},\lambda_{2}\}$. First, we train the model for $10^{5}$ epochs using Adam optimizer Kingma and Ba (2014) with learning rate of $\mu=10^{-3}$. To reduce computation time, we have implemented a mixed precision technique Micikevicius et al. (2017). Then we use L-BFGS-B optimizer Byrd et al. (1995) until the tolerance reaches machine precision. The training is carried out on a single NVIDIA A100 GPU. The code is written using the PyTorch library Paszke et al. (2019) and is openly available at [rep, ].
Figure 4: Velocity (left), pressure (middle) and force (right) probability density functions for DNS data and PINN models trained with various noise levels in the input data. All figures share the same legend. The dashed line corresponds to expression (11).
IV Results
@@
To characterize the statistical properties of reconstructed dense velocity, pressure and force fields, we calculate the corresponding probability density functions (PDFs); see Fig. 4. To illustrate, we used trained models with pressure and force errors close to the mean values presented in Table 2. The velocity PDFs are almost identical to the DNS data even for PINN models based on noisy input data, confirming the low $L_{2}$ error between them. The pressure field is reconstructed worse, and the main difference from the DNS data lies in the typical values, and not in the tails of the distribution function. The force PDF corresponding to expression (2) can be calculated analytically
$\rho(f)=\dfrac{2}{\pi f_{0}}\left(1-\dfrac{f^{2}}{f_{0}^{2}}\right)^{-1/2},$ (11)
where $f=(f_{x}^{2}+f_{y}^{2})^{1/2}\leq f_{0}$, and it diverges at $f\to f_{0}$. The largest disparities between the reconstructed PDFs and the DNS data are concentrated around this point.
Figure 5: Energy spectra for DNS data and PINN models trained with various noise levels in the input data.
In Fig. 5, we compare the energy spectra of the reference flow and PINN reconstructions, which are commonly used to characterize the statistical properties of a velocity field. Since the observation region is not periodic, when constructing the spectra we applied a Hann window function. For this reason, the DNS spectrum is slightly different from that presented earlier in Fig. 2b. As expected from the previous analysis, the overall errors are small. They are mainly concentrated on small scales with $k\gg k_{f}$ and gradually rise with increasing noise level in the input data.
V Conclusion
We showed that PINN can combine measurement data with an accurate physical model to extract additional information from ordinary flow observations. The developed technique can be built on top of current PIV/PTV workflows, potentially improving their functionality. In particular, we investigated a forced two-dimensional turbulent fluid flow and demonstrated that the proposed method can reconstruct the dense velocity and pressure fields, infer the driving force, and determine the unknown fluid viscosity and bottom friction coefficient using only sparse and noisy velocity measurements.
The suggested technique trains two independent NNs simultaneously, one to estimate the velocity and pressure field and the other to approximate the external forcing. This enables us to use different designs in NNs, which makes it possible to take into account the physical features of the considered system. In our case, this allowed us to take into account the stationarity of the external forcing implemented in the physical model. Training of NNs involves minimization of the loss function that penalizes deviations from measured data and violations of the Navier-Stokes equation and incompressibility conditions. While our intention was to select the optimal NN design and hyperparameter configuration, a thorough examination of this issue was outside the scope of this work. We anticipate that the introduction of more advanced NN architectures Zhao, Ding, and Prakash (2023); Wang et al. (2024); Liu et al. (2024); Buzaev et al. (2023) can further improve the developed method.
The reconstructed flows were compared to the reference flow in terms of $L_{2}$ errors, PDFs of the reconstructed fields, and energy spectra. Overall, the results show good accuracy for super-resolution and inference applications and demonstrate moderate robustness to noise in the input data. At low noise levels, the algorithm can correct errors in the velocity field measurements based on information from the physical model. The developed technique can be useful for analyzing flow observations when the physical model of the phenomenon is well constructed.
@@ Acknowledgements.
Data AVAILABILITY
The data that support the findings of this study are openly available at https://github.com/parfenyev/2d-turb-PINN/.